The typical consumer’s food basket in the base year 2015 is as follows:

        30 chickens at $4 each

        10 hams at $5 each

        10 steaks at $8 each

A chicken feed shortage causes the price of chickens to rise to $5.00 each in the year 2016. Hams rise to $7.00 each, and the price of steaks is unchanged.

a. Calculate the change in the “cost-of-eating” index between 2015 and 2016.
  

Year Cost of the basket
2015 $  
2016 $  


Instructions: Enter your responses rounded to one decimal place.

The official cost-of-eating index has   (Click to select)   increased   decreased  by  %.


b. Suppose that consumers are completely indifferent between two chickens and one ham. For this example, how large is the substitution bias in the official “cost-of-eating” index?

The   (Click to select)   increase   decrease  in the cost-of-eating index is  %. The   (Click to select)   underestimate   overestimate  of inflation in the cost of eating reflects substitution bias.
Answer #1

Answer:

Price

2015($)

2016

30 Chicken

$4 each

$5 each

10 hams

$5 each

$7 each

10 steaks

$8 each

$8 each

A]

Calculate cost of eating index between 2015 and 2016 as follows:

Cost of basket(2015) = (30 * 4) + (10*5) + (10*8) = 120 + 50 + 80 = 250

Cost of basket(2016) = (30*5) + (10*7) + (10*8) = 150 + 70 + 80 = 300

The % of cost increase

% increase = 300 – 250/ 250 * 100

= 20%

Thus percentage increase is 20%

B]

Suppose a consumer is indifferent between 2 chickens and one ham. Then people would buy 15 hams for 30 chickens as it is less expensive. The cost of basket in 2016 will be:

Cost of basket(2016) = (7*(10+15)) + (8*10)

= (7*25) + 80

= 255

%increase = 255 – 250 / 250 * 100

= 2%

Thus the percentage increase is 2%